# A Survey of Artificial Intelligence in Gait-Based Neurodegenerative Disease Diagnosis

### v1

@@ 4.5. Alzheimer’s Disease (AD)
In (costa2016application, ) (ID = A-1), the authors explore the use of machine learning classifiers including SVM, MLP, RBNs, and DBSs for diagnosing AD based on postural control kinematics. Four classifiers are compared, and the accuracy of AD diagnosis ranges from 91% to 96.6% when combining postural kinematics and Montreal Cognitive Assessment (MoCA) variable. This study suggests that machine learning models can aid in computer-aided diagnosis of AD using postural control kinematics. The study of (ghoraani2021detection, ) (ID = A-3) collects walking data and derives gait features from participants by using both the ”single-tasking” and ”dual-tasking’ test. The features are then fed as input to the SVM model to classify HC and AD, achieving 78.0% accuracy. In (seifallahi2022alzheimer, ) (ID = A-4), the authors explore the feasibility of using the Timed Up and Go (TUG) test, a simple balance and walking assessment, as a tool for detecting AD compared to HC. Joint position data was collected from subjects performing the TUG test using a Kinect camera. Through signal processing and statistical analyses, significant features were identified and used with a support vector machine classifier, achieving high accuracy (97.75%). Different ML models (SVM, KNN, RF, MLP, NB, MLP) (jeon2023early, ) (ID = A-5) are assembled to diagnose AD using gait parameters extracted from IMU sensors signals, and achieves about 72.2% accuracy. This study totally proposes seven walking experiment paradigms and collects data using seven wearable devices.
4.6. Amyotrophic Lateral Sclerosis (ALS)
@@
In (xia2016novel, ) (Table 5, ID = S-1), the study focuses on analyzing gait variability in patients with amyotrophic lateral sclerosis (ALS) and extracting two key features, standard deviation statistics and permutation entropy. These features were inputted into a support vector machine classifier to classify ALS patients and healthy controls. The results demonstrated high accuracy (92.86%) in distinguishing ALS patients from controls, indicating the potential of gait variability analysis as a diagnostic or monitoring tool for ALS. (felix2020effective, ) (ID = S-2) leverages one-minute gait series and parameters to classify between ALS patients and healthy subjects using machine learning models (SVM, KNN, DT) and achieves up to 96.6% accuracy with DT. (felix2021automatic, ) (ID = S-3) extracts metrics of fluctuation magnitude and fluctuation dynamics from gait time series, and leverages five classic machine learning models (SVM, KNN, NB, LDA, DT) to classify ALS and healthy subjects. Both SVM and KNN achieves the best performance with 89.7% accuracy on the GaitNDD database.
5. Discussions
@@
5.1. Challenges and Future Directions
5.1.1. Limited NDs Types and Studies
@@
5.1.2. Scarcity and Imbalance of Data
The medical data including gait data collected for NDs diagnosis are highly limited, and in the selected 164 articles only several datasets are publicly available: The Gait in Neurodegenerative Disease Database and Gait in Parkinson’s Disease from Physionet database (goldberger2000physiobank, ) are two most frequently-used used datasets; Other datasets include SDUGait (wang2016gait, ), Floodlight PoC dataset (midaglia2019adherence, ), UK Biobank Dataset (doherty2017large, ), Ga, Ju, Si Groups (frenkel2005treadmill, ; hausdorff2007rhythmic, ; yogev2005dual, ). However, these datasets are limited in both sample sizes and NDs types. Most of them contain less than 100 subjects (see Sec. 3.2.6), and the number of NDs types is typically less than 3, while PD gait data are much more than other NDs in terms of quantity and quality ($i.e.,$ unbalanced data size of different diseases). This could limit the learning performance of AI models especially deep learning models, which inherently require large-scale and class-balanced data to achieve more reliable prediction.
To address these challenges, more high-quality gait data of different NDs should be collected. First, it is feasible to cooperate with larger hospitals, medical institutes, or communities to involve more patients and health controls into the collection of gait and related data. Second, the number of patients in each disease and the data size of each patient should be controlled within a reasonable range, $e.g.,$ enlarging the sample size of each disease to more than hundreds and keeping the total data size of each disease as similar as possible. Such balanced data could benefit the model to learn higher inter-class difference with better robustness against intra-class diversity (johnson2019survey, ). To address the challenges of data scarcity and imbalance in this area and facilitate related research, we will collect and open a large-scale multi-modal gait dataset for NDs diagnosis in the future.
5.1.3. Lack of Multi-Source/Modal Gait Data
@@
Therefore, an important future direction is to collect gait data from multiple sources and modalities. First, it is feasible to combine multiple devices as different sources of same-type data ($e.g.,$ multi-view data with differently-positioned cameras), where all devices/sources can compensate each other to better reconstruct or depict the whole pattern with less information loss (liang2020behavioral, ). Second, combining different modalities such as RGB images, depth images, 3D skeletons, and force data (vGRF) is valuable as they can provide gait information from different dimensions ($e.g.,$ appearances, poses) to better predict NDs (zhao2021multimodal, ).
5.1.4. Challenges in AI Model Design
@@
(1) Accuracy: Existing studies report the model accuracy typically ranging from 60% to around 95%, with very few studies possessing accuracy less than 60%. Many studies using C-ML models report model performance with very high accuracy. For the low-accuracy methods, they cannot reach the requirement to assist in medical diagnosis, as the frequent wrong prediction might induce serious medical risk. However, for the methods with very high accuracy, careful evaluations with other indicators such as generalizability and interpretability are required. It is worth noting that these models with nearly 100% accuracy possibly overfit the training data while the the testing data may possess highly similar data distribution to training data ($i.e.,$ small domain shift (sun2016return, )). In this case, these AI models are hard to get satisfactory performance on out-of-distribution data such as new real-world gait data.
@@
(2) Generalizability: Despite high accuracy of quite a few AI methods in this area, many of them are trained, validated, and tested on a single dataset or self-collected gait data, which usually contain limited data sizes, views, scenes or conditions. Therefore, these methods may only perform well on the scenarios that are similar to that of the training data, while they cannot generalize to more challenging data (bar2020impact, ). For example, if the training data contain only gait data of elderly patients, the trained AI model could possess lower accuracy when predicting NDs of young adults using corresponding gait data. In summary, the AI model trained on a single dataset or datasets with limited scenarios ($e.g.,$ single setting) typically possesses low generalization ability in real-world application.
@@
(3) Efficiency: As shown in Sec. 3.2.5, the majority of existing studies (104 papers, 63.4%) utilize C-ML models, which typically enjoy much fewer model parameters and less computational complexity than deep learning models. In general, C-ML models is more efficient than DL models (C-DL and A-DL) under the same accuracy, which makes C-ML more favorable to be deployed in real-world applications. However, compared with conventional ML models, the emerging DL models have stronger ability in mining effective patterns from large-scale medical data, which are more useful and popular in the big data era. For DL models, it is necessary to achieve a good trade-off between model efficiency ($e.g.$, model parameter size, computational complexity) and model performance ($e.g.$, accuracy, speed) (menghani2023efficient, ), so as to better assist in disease diagnosis.
@@
(4) Interpretability: In high-stake fields like medical diagnosis, it is important for AI models to provide human-friendly reliable explanations to facilitate the medical decision (tjoa2020survey, ; payrovnaziri2020explainable, ; zhang2022applications, ). In the surveyed area, there are only a few C-ML model based studies that have identified the important body parts or gait features in classification, while most DL models do not devise an effective mechanism such as feature disentangling and importance visualization (zhang2018visual, ) to provide interpretability. This may increase the risk of wrong diagnosis and discourage the users such as doctors from adopting the predicted result.
@@
To address the aforementioned challenges, it is essential to improve not only the accuracy of an AI model but also its generalizability, efficiency, and interpretability. A potential future direction is to exploit more advanced deep learning models and larger-scale gait datasets with different scenario settings to learn both domain-specific ($e.g.,$ AD-specific) and domain-general ($i.e.,$ NDs-shared) gait features, so as to improve the model accuracy and enable it to be transferred/generalized to other scenarios. For efficiency, it is feasible to employ lighter network architecture ($e.g.,$ MobileNets (howard2017mobilenets, )) or model compression techniques ($e.g.,$ knowledge distillation (gou2021knowledge, ), network pruning (liu2018rethinking, )) to reduce the model size and computational complexity but retaining similar model accuracy. Another promising direction is to adopt smaller gait data such as the emerging 3D skeleton data, which can concisely represent human poses with key body joints and typically require lightweight AI models to learn discriminative gait patterns (rao2021self, ). To improve model interpretability, different types of human-friendly explanation including text description and feature visualization can be considered. We can integrate model-specific explanation mechanisms into different AI architectures, $e.g.,$ visual class activation maps (CAM) (zhou2016learning, ) for CNN, knowledge graphs (ji2021survey, ) for graph neural networks (GNN). Inspired by the recent success of ChatGPT, it can be used to provide insightful description for the importance of features if we transform gait features into text format. Moreover, ChatGPT can also act as an agent to lead and improve the AI model learning (haupt2023ai, ), in which we can add prompts to generate interpretable results. As there exist key body parts and poses for NDs diagnosis, we can try to disentangle the gait abnormality related semantics from normal motion semantics ($e.g.,$ daily actions) to identify the most important or unique component of NDs patients’ features. Simultaneously improving the accuracy, generalizability, efficiency, and interpretability, it is hopefully to obtain a more reliable and practical AI model for gait-based NDs diagnosis. Meanwhile, a unified evaluation protocol, which synergizes different benchmark datasets, validation settings, and evaluation metrics, should be further devised to fairly compare different AI models to facilitate the development of this area.
@@
Figure 9. Overview of 3D skeleton based NDs diagnosis framework. We visualize different directions of AI model design in terms of skeleton-level ((1)-(3)) and sequence-level NDs diagnosis ((a)-(c)), and also provide representative AI methods of this direction as an example solution (${[1]}$: (rao2023hierarchical, ); ${[2]}$: (rao2021sm, ); ${[3]}$: (andersson2015person, ); ${[4]}$: (rao2023transg, )); ${[5]}$: (rao2022simmc, ))
5.2. Research Vision
@@
Recently driven by economical, non-obtrusive and accurate skeleton-tracking devices like Kinect (shotton2011real-time, ), 3D skeleton data has been a popular and generic data modality for many gait-related tasks such as gait recognition and person re-identification (pala2019enhanced, ; liao2020model, ; andersson2015person, ; munaro20143d, ; munaro2014one, ; barbosa2012re, ; rao2020self, ; rao2021self, ; rao2021sm, ; rao2021multi, ). A 3D skeleton is defined as 3D coordinates of key human joints (typically 20 or 25 key joints (shotton2011real-time, )) of a person, while 3D skeletons are defined as all joints’ temporal series conveying motion dynamics of the person. In this concise way, 3D skeletons can provide both body structure and pose information for gait analysis. Encouraged by the promising results of 3D skeletons in many fields, we propose the research vision of exploiting 3D skeleton data to characterize NDs-related gait patterns to enhance the diagnosis performance. In this section, we first discuss the unique advantages of 3D skeleton data for gait representation and learning, which can potentially address challenges in AI-based NDs diagnosis in terms of efficiency and scalability (see Sec. 5.2.1). Then, we propose a generic 3D skeleton based framework to design AI models specifically for NDs diagnosis (see Sec. 5.2.2).
5.2.1. Skeletons as Efficient and Scalable Gait Representations
@@
Specifically, 3D skeleton data possess the following advantages to improve the efficiency of gait learning: (a) Smaller data inputs compared with appearance-based methods that typically require large-size ($e.g.,$ high-resolution) image/video data; (b) Lighter AI models than conventional image-based models to process and learn skeletons (rao2022simmc, ); (c) Lower resource/device requirement compared with other sensor-based data that require multiple devices such as IMUs or expensive systems such as walkway systems; (d) Higher convenience of collection with unobtrusive and contactless detection compared with other sensor methods that usually require wearing devices ($e.g.,$ sensors) or body markers. These advantages enable them to be potentially applied to capturing and learning gait with higher efficiency in terms of real-world deployment and model design.
@@
On the other hand, 3D skeletons can serve as scalable gait representations for different application scenarios: (a) For medical diagnosis and other privacy-sensitive areas, 3D skeletons can replace traditional vision data such as RGB images to exclude all visual appearance information ($e.g.,$ faces of patients), so as to provide better privacy protection for downstream tasks such as gait classification; (b) Compared with other gait representations that rely on RGB or depth images, 3D skeleton based representations possess more robust performance under variations of scales, views or other external factors (han2017space, ), which enables them to be flexibly used in more scenarios; (c) As 3D skeletons are general modality for many tasks including action recognition (rao2021augmented, ), gait recognition (liao2020model, ), and person re-identification (rao2021self, ), all these tasks could be flexibly transferred or combined to provide better abnormal gait classification for NDs diagnosis. For example, the most distinctive pose regions in action recognition tasks can be utilized to locate the gait abnormality of NDs patients, while the identity-specific gait patterns in person re-identification tasks can be exploited to recognize NDs patients and help decouple disease-specific gait patterns; (d) There are many successful practices to combine 3D skeleton data and other modalities such as RGB images, depth images, and 3D clouds (pala2019enhanced, ; munaro20143d, ) to improve feature effectiveness for various tasks. It is feasible to combine 3D skeleton data with the above data modalities or other data such as IMUs signals and force signals to further boost the performance of gait learning and NDs diagnosis. It is worth noting that the most commonly used skeleton-tracking device, Kinect, can simultaneously capture RGB images, depth images, and 3D skeletons, which provides great benefit for multi-modal representations of gait.
5.2.2. 3D Skeleton based NDs Diagnosis Framework
@@
As presented in Fig. 9, by using 3D skeletons as gait representations, NDs can be diagnosed from two main aspects, a single skeleton and a skeleton sequence. For skeleton-level NDs diagnosis, the abnormalities in body structure, joint relations, and skeletal properties are key focus of AI models: (1) We can exploit multi-level or hierarchical body structure modeling (rao2023hierarchical, ) to help capture more discriminative features between NDs patients and healthy subjects; (2) The structural and actional relations between body joints or parts usually characterize unique walking patterns of a person (murray1964walking, ), so we can combine them to learn both local and global motion correlations to better catch valuable patterns (rao2021sm, ) ($e.g.,$ abnormal gait patterns); (3) There exist crucial skeletal properties including anthropometric and geometric attributes ($e.g.,$ joint distances and angles) that can help recognize different actions (rao2021augmented, ) or identify different persons (andersson2015person, ). These key skeletal properties incorporating domain-specific knowledge could be potentially used to identify different NDs.
@@
For skeleton sequence level NDs diagnosis, learning NDs-related pose dynamics, gait attributes, and latent motion semantics is the key to AI model design (see Fig. 9): (a) For NDs patients, their poses might be unnatural or twisted, which could be reflected on their motion dynamics such as body-joint trajectory. Therefore, it is feasible to capture pose or trajectory dynamics from 3D skeletons (rao2023transg, ) and extract the NDs-related parts for diagnosis; (b) Considering that NDs patients typically possess abnormal gait attributes such as lower speed and uncertain stride length, we can compute these gait features from 3D skeletons (andersson2015person, ) as important discriminators of different diseases; (c) Advanced AI models such as deep learning models can learn latent high-level semantics in high-dimensional feature space to help better classify abnormal gaits. For example, we may devise self-supervised tasks such as skeleton sequence reconstruction, sorting, prediction (rao2021self, ), or contrasting (rao2022simmc, ) to learn high-level semantics in terms of motion continuity and consistency, which can be used to detect gait abnormality by comparing with normal samples for NDs diagnosis. We can also leverage unsupervised clustering algorithms to find latent gait semantics ($e.g.,$ gait prototypes (rao2022simmc, )) related to different NDs for diagnosis.
6. Conclusion

### v2

@@ 4.6 Alzheimer’s Disease (AD)
In [221] (ID = A-1), the authors explore the use of machine learning classifiers including SVM, MLP, RBNs, and DBSs for diagnosing AD based on postural control kinematics. Four classifiers are compared, and the accuracy of AD diagnosis ranges from 91% to 96.6% when combining postural kinematics and Montreal Cognitive Assessment (MoCA) variable. This study suggests that machine learning models can aid in computer-aided diagnosis of AD using postural control kinematics. [71] (ID = A-2) proposes a cascade neural network approach that combines gait and EEG data for more accurate and efficient classification of AD. The model utilizes attention-based spatial temporal graph convolutional networks (ST-GCN) to extract discriminative features from gait data. Experimental results show superior performance compared to other AD diagnosis methods, with the lower body and right upper limb identified as important for early AD diagnosis. The study of [222] (ID = A-3) collects walking data and derives gait features from participants by using both the ”single-tasking” and ”dual-tasking’ test. The features are then fed as input to the SVM model to classify HC and AD, achieving 78.0% accuracy. In [107] (ID = A-4), the authors explore the feasibility of using the Timed Up and Go (TUG) test, a simple balance and walking assessment, as a tool for detecting AD compared to HC. Joint position data was collected from subjects performing the TUG test using a Kinect camera. Through signal processing and statistical analyses, significant features were identified and used with a support vector machine classifier, achieving high accuracy (97.75%). Different ML models (SVM, KNN, RF, MLP, NB, MLP) [223] (ID = A-5) are assembled to diagnose AD using gait parameters extracted from IMU sensors signals, and achieves about 72.2% accuracy. This study totally proposes seven walking experiment paradigms and collects data using seven wearable devices.
4.7 Amyotrophic Lateral Sclerosis (ALS)
@@
In [104] (Table 7, ID = S-1), the study focuses on analyzing gait variability in patients with amyotrophic lateral sclerosis (ALS) and extracting two key features, standard deviation statistics and permutation entropy. These features were inputted into a support vector machine classifier to classify ALS patients and healthy controls. The results demonstrated high accuracy (92.86%) in distinguishing ALS patients from controls, indicating the potential of gait variability analysis as a diagnostic or monitoring tool for ALS. [224] (ID = S-2) leverages one-minute gait series and parameters to classify between ALS patients and healthy subjects using machine learning models (SVM, KNN, DT) and achieves up to 96.6% accuracy with DT. [225] (ID = S-3) extracts metrics of fluctuation magnitude and fluctuation dynamics from gait time series, and leverages five classic machine learning models (SVM, KNN, NB, LDA, DT) to classify ALS and healthy subjects. Both SVM and KNN achieves the best performance with 89.7% accuracy on the GaitNDD database.
5 Current Challenges and Future Directions
@@
5.1 Limited NDs Types and Studies
@@
5.2 Scarcity and Imbalance of Data
The medical data including gait data collected for NDs diagnosis are highly limited, and in the selected 164 articles only several datasets are publicly available: The Gait in Neurodegenerative Disease Database and Gait in Parkinson’s Disease from Physionet database [234] are two most frequently-used used datasets; Other datasets include SDUGait [235], Floodlight PoC dataset [236], UK Biobank Dataset [237], Ga, Ju, Si Groups [238, 239, 240]. However, these datasets are limited in both sample sizes and NDs types. Most of them contain less than 100 subjects (see Sec. 3.2.6), and the number of NDs types is typically less than 3, while PD gait data are much more than other NDs in terms of quantity and quality ($i.e.,$ unbalanced data size of different diseases). This could limit the learning performance of AI models especially deep learning models, which inherently require large-scale and class-balanced data to achieve more reliable prediction.
To address these challenges, more high-quality gait data of different NDs should be collected. First, it is feasible to cooperate with larger hospitals, medical institutes, or communities to involve more patients and health controls into the collection of gait and related data. Second, the number of patients in each disease and the data size of each patient should be controlled within a reasonable range, $e.g.,$ enlarging the sample size of each disease to more than hundreds and keeping the total data size of each disease as similar as possible. Such balanced data could benefit the model to learn higher inter-class difference with better robustness against intra-class diversity [241]. To address the challenges of data scarcity and imbalance in this area and facilitate related research, we will collect and open a large-scale multi-modal gait dataset for NDs diagnosis in the future.
5.3 Lack of Multi-Source/Modal Gait Data
@@
Therefore, an important future direction is to collect gait data from multiple sources and modalities. First, it is feasible to combine multiple devices as different sources of same-type data ($e.g.,$ multi-view data with differently-positioned cameras), where all devices/sources can compensate each other to better reconstruct or depict the whole pattern with less information loss [242]. Second, combining different modalities such as RGB images, depth images, 3D skeletons, and force data (vGRF) is valuable as they can provide gait information from different dimensions ($e.g.,$ appearances, poses) to better predict NDs [77].
5.4 Challenges in AI Model Design
@@
(1) Accuracy: Existing studies report the model accuracy typically ranging from 60% to around 95%, with very few studies possessing accuracy less than 60%. Many studies using C-ML models report model performance with very high accuracy. For the low-accuracy methods, they cannot reach the requirement to assist in medical diagnosis, as the frequent wrong prediction might induce serious medical risk. However, for the methods with very high accuracy, careful evaluations with other indicators such as generalizability and interpretability are required. It is worth noting that these models with nearly 100% accuracy possibly overfit the training data while the the testing data may possess highly similar data distribution to training data ($i.e.,$ small domain shift [243]). In this case, these AI models are hard to get satisfactory performance on out-of-distribution data such as new real-world gait data.
@@
(2) Generalizability: Despite high accuracy of quite a few AI methods in this area, many of them are trained, validated, and tested on a single dataset or self-collected gait data, which usually contain limited data sizes, views, scenes or conditions. Therefore, these methods may only perform well on the scenarios that are similar to that of the training data, while they cannot generalize to more challenging data [244]. For example, if the training data contain only gait data of elderly patients, the trained AI model could possess lower accuracy when predicting NDs of young adults using corresponding gait data. In summary, the AI model trained on a single dataset or datasets with limited scenarios ($e.g.,$ single setting) typically possesses low generalization ability in real-world application.
@@
(3) Efficiency: As shown in Sec. 3.2.5 and Table 1, the majority of existing studies (104 papers, 63.4%) utilize C-ML models, which typically enjoy much fewer model parameters and less computational complexity than deep learning models. In general, C-ML models is more efficient than DL models (C-DL and A-DL) under the same accuracy, which makes C-ML more favorable to be deployed in real-world applications. However, compared with conventional ML models, the emerging DL models have stronger ability in mining effective patterns from large-scale medical data, which are more useful and popular in the big data era. For DL models, it is necessary to achieve a good trade-off between model efficiency ($e.g.$, model parameter size, computational complexity) and model performance ($e.g.$, accuracy, speed) [245], so as to better assist in disease diagnosis.
@@
(4) Interpretability: In high-stake fields like medical diagnosis, it is important for AI models to provide human-friendly reliable explanations to facilitate the medical decision [246, 247, 248]. In the surveyed area, there are only a few C-ML model based studies that have identified the important body parts or gait features in classification, while most DL models do not devise an effective mechanism such as feature disentangling and importance visualization [249] to provide interpretability. This may increase the risk of wrong diagnosis and discourage the users such as doctors from adopting the predicted result.
@@
To address the aforementioned challenges, it is essential to improve not only the accuracy of an AI model but also its generalizability, efficiency, and interpretability. A potential future direction is to exploit more advanced deep learning models and larger-scale gait datasets with different scenario settings to learn both domain-specific ($e.g.,$ AD-specific) and domain-general ($i.e.,$ NDs-shared) gait features, so as to improve the model accuracy and enable it to be transferred/generalized to other scenarios. For efficiency, it is feasible to employ lighter network architecture ($e.g.,$ MobileNets [250]) or model compression techniques ($e.g.,$ knowledge distillation [251], network pruning [252]) to reduce the model size and computational complexity but retaining similar model accuracy. Another promising direction is to adopt smaller gait data such as the emerging 3D skeleton data, which can concisely represent human poses with key body joints and typically require lightweight AI models to learn discriminative gait patterns [253]. To improve model interpretability, different types of human-friendly explanation including text description and feature visualization can be considered. We can integrate model-specific explanation mechanisms into different AI architectures, $e.g.,$ visual class activation maps (CAM) [254] for CNN, knowledge graphs [255] for graph neural networks (GNN). Inspired by the recent success of ChatGPT, it can be used to provide insightful description for the importance of features if we transform gait features into text format. Moreover, ChatGPT can also act as an agent to lead and improve the AI model learning [256], in which we can add prompts to generate interpretable results. As there exist key body parts and poses for NDs diagnosis, we can try to disentangle the gait abnormality related semantics from normal motion semantics ($e.g.,$ daily actions) to identify the most important or unique component of NDs patients’ features. Simultaneously improving the accuracy, generalizability, efficiency, and interpretability, it is hopefully to obtain a more reliable and practical AI model for gait-based NDs diagnosis. Meanwhile, a unified evaluation protocol, which synergizes different benchmark datasets, validation settings, and evaluation metrics, should be further devised to fairly compare different AI models to facilitate the development of this area.
@@
Figure 9: Overview of 3D skeleton based NDs diagnosis framework. We visualize different directions of AI model design in terms of skeleton-level ((1)-(3)) and sequence-level NDs diagnosis ((a)-(c)), and also provide representative AI methods of this direction as an example solution (${[1]}$: [257]; ${[2]}$: [258]; ${[3]}$: [259]; ${[4]}$: [260]); ${[5]}$: [261])
5.5 Research Vision: Application of 3D Skeleton Data
@@
Recently driven by economical, non-obtrusive and accurate skeleton-tracking devices like Kinect [262], 3D skeleton data has been a popular and generic data modality for many gait-related tasks such as gait recognition and person re-identification [263, 264, 259, 265, 266, 267, 268, 253, 258, 269]. A 3D skeleton is defined as 3D coordinates of key human joints (typically 20 or 25 key joints [262]) of a person, while 3D skeletons are defined as all joints’ temporal series conveying motion dynamics of the person. In this concise way, 3D skeletons can provide both body structure and pose information for gait analysis. Encouraged by the promising results of 3D skeletons in many fields, we propose the research vision of exploiting 3D skeleton data to characterize NDs-related gait patterns to enhance the diagnosis performance. In this section, we first discuss the unique advantages of 3D skeleton data for gait representation and learning, which can potentially address challenges in AI-based NDs diagnosis in terms of efficiency and scalability (see Sec. 5.5.1). Then, we propose a generic 3D skeleton based framework to design AI models specifically for NDs diagnosis (see Sec. 5.5.2).
5.5.1 Skeletons as Efficient and Scalable Gait Representations
@@
Specifically, 3D skeleton data possess the following advantages to improve the efficiency of gait learning: (a) Smaller data inputs compared with appearance-based methods that typically require large-size ($e.g.,$ high-resolution) image/video data; (b) Lighter AI models than conventional image-based models to process and learn skeletons [261]; (c) Lower resource/device requirement compared with other sensor-based data that require multiple devices such as IMUs or expensive systems such as walkway systems; (d) Higher convenience of collection with unobtrusive and contactless detection compared with other sensor methods that usually require wearing devices ($e.g.,$ sensors) or body markers. These advantages enable them to be potentially applied to capturing and learning gait with higher efficiency in terms of real-world deployment and model design.
@@
On the other hand, 3D skeletons can serve as scalable gait representations for different application scenarios: (a) For medical diagnosis and other privacy-sensitive areas, 3D skeletons can replace traditional vision data such as RGB images to exclude all visual appearance information ($e.g.,$ faces of patients), so as to provide better privacy protection for downstream tasks such as gait classification; (b) Compared with other gait representations that rely on RGB or depth images, 3D skeleton based representations possess more robust performance under variations of scales, views or other external factors [270], which enables them to be flexibly used in more scenarios; (c) As 3D skeletons are general modality for many tasks including action recognition [271], gait recognition [264], and person re-identification [253], all these tasks could be flexibly transferred or combined to provide better abnormal gait classification for NDs diagnosis. For example, the most distinctive pose regions in action recognition tasks can be utilized to locate the gait abnormality of NDs patients, while the identity-specific gait patterns in person re-identification tasks can be exploited to recognize NDs patients and help decouple disease-specific gait patterns; (d) There are many successful practices to combine 3D skeleton data and other modalities such as RGB images, depth images, and 3D clouds [263, 265] to improve feature effectiveness for various tasks. It is feasible to combine 3D skeleton data with the above data modalities or other data such as IMUs signals and force signals to further boost the performance of gait learning and NDs diagnosis. It is worth noting that the most commonly used skeleton-tracking device, Kinect, can simultaneously capture RGB images, depth images, and 3D skeletons, which provides great benefit for multi-modal representations of gait.
5.5.2 3D Skeleton based NDs Diagnosis Framework
@@
As presented in Fig. 9, by using 3D skeletons as gait representations, NDs can be diagnosed from two main aspects, a single skeleton and a skeleton sequence. For skeleton-level NDs diagnosis, the abnormalities in body structure, joint relations, and skeletal properties are key focus of AI models: (1) We can exploit multi-level or hierarchical body structure modeling [257] to help capture more discriminative features between NDs patients and healthy subjects; (2) The structural and actional relations between body joints or parts usually characterize unique walking patterns of a person [272], so we can combine them to learn both local and global motion correlations using AI models ($e.g.,$ graph attention networks [273]) to better catch valuable patterns [258] ($e.g.,$ abnormal gait patterns); (3) There exist crucial skeletal properties including anthropometric and geometric attributes ($e.g.,$ joint distances and angles) that can help recognize different actions [271] or identify different persons [259]. These key skeletal properties incorporating domain-specific knowledge could be potentially used to identify different NDs.
@@
For skeleton sequence level NDs diagnosis, learning NDs-related pose dynamics, gait attributes, and latent motion semantics is the key to AI model design (see Fig. 9): (a) For NDs patients, their poses might be unnatural or twisted, which could be reflected on their motion dynamics such as body-joint trajectory. Therefore, it is feasible to capture pose or trajectory dynamics from 3D skeletons [260] and extract the NDs-related parts for diagnosis; (b) Considering that NDs patients typically possess abnormal gait attributes such as lower speed and uncertain stride length, we can compute these gait features from 3D skeletons [259] as important discriminators of different diseases; (c) Advanced AI models such as deep learning models can learn latent high-level semantics in high-dimensional feature space to help better classify abnormal gaits. For example, we may devise self-supervised tasks such as skeleton sequence reconstruction, sorting [253], forecasting [274], or contrasting [261] to learn high-level semantics in terms of motion continuity and consistency, which can be used to detect gait abnormality by comparing with normal samples for NDs diagnosis. We can also leverage unsupervised clustering algorithms to find latent gait semantics ($e.g.,$ gait prototypes [261]) related to different NDs for diagnosis.
6 Conclusion